# On the numerical approximation of hyperbolic mean curvature flows for surfaces

### v1

@@ 1 Introduction
for a family of closed oriented surfaces $(\mathcal{S}(t))_{t\in[0,T]}$ in ${\mathbb{R}}^{3}$. Here $\mathcal{V}_{\mathcal{S}}$ denotes the velocity of $(\mathcal{S}(t))_{t\in[0,T]}$ in the direction of the unit normal $\nu_{\mathcal{S}}$, $\partial_{t}^{\text{\tiny$\square$}}$ is the normal time derivative on $(\mathcal{S}(t))_{t\in[0,T]}$, and $H$ denotes the mean curvature of $\mathcal{S}(t)$ with the sign convention that the unit sphere with outward normal satisfies $H=-2$. As for the coefficient function $g:{\mathbb{R}}_{\geq 0}\to{\mathbb{R}}_{\geq 0}$, we shall focus on two choices: Firstly, the choice $g(s)=1$ leads to the equation
@@
which has been derived by LeFloch and Smoczyk in [18] with the help of the Hamiltonian principle. As a result, solutions of (1.3) conserve the sum $E(t)=\tfrac{1}{2}\int_{\mathcal{S}(t)}(\mathcal{V}_{\mathcal{S}}^{2}+2)\;{\rm d}{\mathcal{H}}^{2}$ of kinetic and internal energy, see Corollary 4.7 in [18]. Here, ${\rm d}\mathcal{H}^{2}$ denotes integration with respect to the two-dimensional Hausdorff measure. Recalling [6, Theorem 32], we can derive this property from (1.3) via
@@
The aim of this paper is to develop numerical methods for the evolution of surfaces either by (1.2) or by (1.3). Our primary approach is based on a parametric description of the moving surfaces and discretizes a system of PDEs satisfied by the position vector of the surface. The corresponding systems for both evolution laws are introduced in Section 2, see Lemma 2.1. The systems are chosen in such a way that the resulting flow is normal, i.e. the velocity vector has no tangential component. This property has proven to be helpful both in analytical investigations, see [16, 18], and in the numerical analysis carried out in [11]. Our numerical schemes are introduced in Section 3. We shall adapt a finite element method due to Dziuk, [13], for mean curvature flow, which approximates the evolving surface by a triangulated surface and parameterizes the solution in the next time step over the current surface. Apart from this general case, we are also interested in the more special situation that the evolving surfaces are rotationally symmetric. Here, the evolution can be described in terms of the profile curve and we are able to extend the finite difference approach developed in [11] to this setting. There are two reasons for considering the case of axisymmetric surfaces: on the one hand, an algorithm adapted to this setting will substantially reduce the computational cost, while on the other hand we can use the corresponding numerical results in order to verify our finite element simulations for the general case. Those simulations will be presented in Section 4 for the two flows (1.2) and (1.3), covering both the general and the axisymmetric case. They show interesting oscillatory behaviour of the surfaces and provide numerical evidence for the onset of singularities in three space dimensions.
@@ 2.1 The general case
where ${\rm T}_{z}S(t)$ denotes the tangent space to $S(t)$ at the point $z\in\mathcal{S}(t)$.
@@ 3.1 The general case
Let us now combine (3.2), (3.3) and (3.4) to define time-discrete approximations for (2.1). In what follows it will be convenient to replace $\nabla_{\mathcal{S}^{m}}{\rm id}$ by $\frac{1}{2}\nabla_{\mathcal{S}^{m}}\bigl{(}X^{m+1}+(X^{m})^{-1}\bigr{)}$ in (3.4) in order to enhance the stability of the resulting schemes.
@@
In the remainder of the paper we will always assume that $V_{0}$ is constant so that the above relation simplifies to
@@
For a finite element discretization we closely follow [13]. Let $\mathcal{S}^{m}_{h}$ be a polyhedral surface, i.e. a union of nondegenerate triangles with no hanging vertices (see [9, p. 164]), approximating the closed surface $\mathcal{S}(t_{m})$. Let $\mathcal{S}^{m}_{h}=\bigcup_{j=1}^{J}\overline{\sigma^{m}_{j}}$, where $\{\sigma^{m}_{j}\}_{j=1}^{J}$ is a family of mutually disjoint open triangles with vertices $\{p^{m}_{k}\}_{k=1}^{K}$ and set $h_{m}:=\max_{j=1\to J}{\rm diam}(\sigma^{m}_{j})$. Then let
@@
while in the case $g(s)=1+\tfrac{1}{2}$
@@ 3.2 The axisymmetric case
$\kappa\nu-\frac{\nu\cdot e_{1}}{x\cdot e_{1}}\nu=\frac{1}{|x_{\rho}|}\bigl{(}\frac{x_{\rho}}{|x_{\rho}|}\bigr{)}_{\rho}-\frac{1}{|x_{\rho}|}\frac{x_{\rho}\cdot e_{2}}{x\cdot e_{1}}\tau^{\perp}\quad\mbox{ in }I\times(0,T],$
@@
$\tfrac{1}{2}(q^{m}_{j}+q^{m}_{j+1})h\frac{x^{m+1}_{j}-2x^{m}_{j}+x^{m-1}_{j}}{(\Delta t)^{2}}=g\Bigl{(}\left|\frac{x^{m}_{j}-x^{m-1}_{j}}{\Delta t}\right|^{2}\Bigr{)}\left[\frac{1}{2q^{m}_{j+1}}\delta^{-}x^{m+1}_{j+1}-\frac{1}{2q^{m}_{j}}\delta^{-}x^{m+1}_{j}\right.$
$\qquad\quad\left.+\frac{1}{2q^{m}_{j+1}}\delta^{-}x^{m-1}_{j+1}-\frac{1}{2q^{m}_{j}}\delta^{-}x^{m-1}_{j}-\frac{\delta^{1}x^{m}_{j}\cdot e_{2}}{x^{m}_{j}\cdot e_{1}}(\theta^{m}_{j})^{\perp}\right]$
@@
$x^{m+1}_{0}\cdot e_{1}=0;\quad\delta^{+}x^{m+1}_{0}\cdot e_{2}=\tfrac{1}{4}h(q^{m}_{1})^{2}\frac{x^{m+1}_{0}-2x^{m}_{0}+x^{m-1}_{0}}{(\Delta t)^{2}}\cdot e_{2},$ (3.16b)
$x^{m+1}_{J}\cdot e_{1}=0;\quad\delta^{-}x^{m+1}_{J}\cdot e_{2}=-\tfrac{1}{4}h(q^{m}_{J})^{2}\frac{x^{m+1}_{J}-2x^{m}_{J}+x^{m-1}_{J}}{(\Delta t)^{2}}\cdot e_{2}.$ (3.16c)
A motivation for the form of the conditions (3.16b), (3.16c) can be found in [10].
@@
$\tfrac{1}{2}(q^{m}_{j}+q^{m}_{j+1})h^{2}\frac{X_{j}}{(\Delta t)^{2}}=\frac{\widehat{g}_{j}}{2q^{m}_{j+1}}(X_{j+1}-X_{j})-\frac{\widehat{g}_{j}}{2q^{m}_{j}}(X_{j}-X_{j-1}),\quad j=1,\ldots,J-1,$ (3.17a)
$X_{0}\cdot e_{1}=0;\quad(X_{1}-X_{0})\cdot e_{2}=\tfrac{1}{4}h^{2}(q^{m}_{1})^{2}\frac{X_{0}}{(\Delta t)^{2}}\cdot e_{2},$ (3.17b)
$X_{J}\cdot e_{1}=0;\quad(X_{J}-X_{J-1})\cdot e_{2}=-\tfrac{1}{4}h^{2}(q^{m}_{J})^{2}\frac{X_{J}}{(\Delta t)^{2}}\cdot e_{2},$ (3.17c)
where $\widehat{g}_{j}=g(|\frac{x^{m}_{j}-x^{m-1}_{j}}{\Delta t}|^{2})$. Clearly, (3.17) decouples into two tridiagonal linear systems for $X\cdot e_{1}$ and $X\cdot e_{2}$, which are both strictly diagonally dominant. Hence it follows that $X=0$.
@@
In addition we have from [4, (2.17)] that
@@ 4 Numerical results
We remark that in all our computations for (1.2) and (1.3), respectively, natural discrete analogues of $\tfrac{1}{2}\int_{\mathcal{S}(t)}(\mathcal{V}_{\mathcal{S}}^{2}+2)\;{\rm d}{\mathcal{H}}^{2}$ and $\int_{\mathcal{S}(t)}e^{\frac{1}{2}\mathcal{V}_{\mathcal{S}}^{2}}\;{\rm d}{\mathcal{H}}^{2}$ were well preserved, recall (1.2) and (1).
@@ 4.1.1 Radial evolutions for a sphere
Our first set of numerical experiments is for the evolution of a sphere. To validate our schemes (3.1) and (3.16) with $g(s)=1$, we begin with a convergence experiment, generalizing the solution from [11, §5.2] to higher space dimensions. It can be shown that a family of spheres in $\mathbb{R}^{3}$ with radius $r(t)$ and outer normal $\nu(t)$ is a solution to (1.2) with $\mathcal{V}_{\mathcal{S}\mid_{t=0}}=V_{0}\in{\mathbb{R}}$ if
@@
and where we employed the time step size choice $\Delta t=0.25h_{0}$. We note that the observed order of convergence approaches the optimal rate $\mathcal{O}(h^{2})$.
@@
and where we have chosen $\Delta t=h$. Also for the scheme (3.16) we obtain an optimal convergence rate.
@@ 4.1.2 Axisymmetric evolutions for nonspherical surfaces
We begin with an experiment for an initial $2:1:1$ cigar, when the initial velocity is chosen as $V_{0}=0$, see Figure 3. Again we observe an excellent agreement between the solutions for our schemes (3.1) and (3.16) with $g(s)=1$. At around $t=0.5$ there seems to be the onset of a singularity. Notice that the first kink in the curvature plot appears to show the flattening of the long part of the surface, which then resolves itself. However, then the shorter part flattens leading to a blow-up of curvature.
@@
A simulation for an initial torus with radii $R=2$ and $r=1$ can be seen in Figure 5 for $V_{0}=0$. Here we see a singularity develop, before the torus can shrink to a circle, when the outer part of the torus becomes flat. This kind of singularity cannot occur for mean curvature flow, and so is unique to the hyperbolic flow considered here. In particular, we recall from [22, Proposition 3] that for mean curvature flow it is known that there exists a critical value $r_{crit}\in(0,1)$, such that for $0<\frac{r}{R}<r_{crit}$ the solution shrinks to a circle, while for $r_{crit}<\frac{r}{R}<1$ it closes up the hole at the origin. This was also confirmed by numerical experiments in e.g. [2, Figs. 5, 6], [5, Figs. 2, 3] and [1, Figs. 2, 3, 4], where the latter also provides an estimate for $r_{crit}$ of about $0.6415$.
@@
We end this section with a simulation for hyperbolic mean curvature flow for a surface that is the boundary of a nonconvex domain. In particular, as the initial discrete surface $\mathcal{S}^{0}_{h}$ we choose the biconcave disk obtained in [3, Fig. 8]. The numerical results for the scheme (3.1) are presented in Figure 7.
@@
In order to qualitatively confirm the evolution shown in Figure 7, we compute a very similar evolution for the axisymmetric finite difference scheme (3.16) with $g(s)=1$. Here we choose as initial data $x^{0}$ the curve displayed in the middle of [7, Fig. 5]. The obtained simulation is shown in Figure 8, which agrees very well with the full 3d results in Figure 7. In particular, at the side of the disk a concavity appears, which then leads to a blow-up in curvature and a resulting singularity for the flow.

### v2

@@ 1 Introduction
for a family of closed oriented surfaces $(\mathcal{S}(t))_{t\in[0,T]}$ in ${\mathbb{R}}^{3}$. Here $\mathcal{V}_{\mathcal{S}}$ denotes the velocity of $(\mathcal{S}(t))_{t\in[0,T]}$ in the direction of the unit normal $\nu_{\mathcal{S}}$, $\partial_{t}^{\text{\tiny$\square$}}$ is the normal time derivative on $(\mathcal{S}(t))_{t\in[0,T]}$, and $H$ denotes the mean curvature of $\mathcal{S}(t)$ with the sign convention that the unit sphere with outward normal satisfies $H=-2$. As for the coefficient function $g:{\mathbb{R}}_{\geq 0}\to{\mathbb{R}}_{>0}$, we shall focus on two choices: Firstly, the choice $g(s)=1$ leads to the equation
@@
which has been derived by LeFloch and Smoczyk in [23] with the help of the Hamiltonian principle. As a result, solutions of (1.3) conserve the sum $E(t)=\tfrac{1}{2}\int_{\mathcal{S}(t)}(\mathcal{V}_{\mathcal{S}}^{2}+2)\;{\rm d}{\mathcal{H}}^{2}$ of kinetic and internal energy, see Corollary 4.7 in [23]. Here, ${\rm d}\mathcal{H}^{2}$ denotes integration with respect to the two-dimensional Hausdorff measure. Recalling [7, Theorem 32], we can derive this property from (1.3) via
@@
The aim of this paper is to develop numerical methods for the evolution of surfaces either by (1.2) or by (1.3). Our primary approach is based on a parametric description of the moving surfaces and discretizes a system of PDEs satisfied by the position vector of the surface. The corresponding systems for both evolution laws are introduced in Section 2, see Lemma 2.1. The systems are chosen in such a way that the resulting flow is normal, i.e. the velocity vector has no tangential component. This property has proven to be helpful both in analytical investigations, see [21, 23], and in the numerical analysis carried out in [12]. We remark that for related parabolic geometric evolution equations, like mean curvature flow, surface diffusion and Willmore flow, several different approaches on introducing suitable tangential motions to help maintain or improve the mesh quality exist, see e.g. [2, 3, 7, 16, 20, 14]. However, for simplicity we will focus on purely normal formulations for the hyperbolic flows considered in this paper. Our numerical schemes are introduced in Section 3. We shall adapt a finite element method due to Dziuk, [15], for mean curvature flow, which approximates the evolving surface by a triangulated surface and parameterizes the solution in the next time step over the current surface. Apart from this general case, we are also interested in the more special situation that the evolving surfaces are rotationally symmetric. Here, the evolution can be described in terms of the profile curve and we are able to extend the finite difference approach developed in [12] to this setting. There are two reasons for considering the case of axisymmetric surfaces: on the one hand, an algorithm adapted to this setting will substantially reduce the computational cost, while on the other hand we can use the corresponding numerical results in order to verify our finite element simulations for the general case. Those simulations will be presented in Section 4 for the two flows (1.2) and (1.3), covering both the general and the axisymmetric case. They show interesting oscillatory behaviour of the surfaces and provide numerical evidence for the onset of singularities in three space dimensions.
@@ 2.1 The general case
where ${\rm T}_{z}\mathcal{S}(t)$ denotes the tangent space to $\mathcal{S}(t)$ at the point $z\in\mathcal{S}(t)$.
@@ 3.1 The general case
Let us now combine (3.2), (3.3) and (3.4) to define time-discrete approximations for (2.1). In what follows it will be convenient to replace $\nabla_{\mathcal{S}^{m}}{\rm id}$ by $\frac{1}{2}\nabla_{\mathcal{S}^{m}}\bigl{(}X^{m+1}+(X^{m})^{-1}\bigr{)}$ in (3.4) in order to enhance the stability and accuracy of the resulting fully discrete schemes. This approach is motivated by the authors’ previous work in [12], as well as by the proposed numerical approximations for the linear wave equation in [18, §2.7].
@@
In the remainder of the paper, to simplify the presention, we will always assume that $V_{0}$ is constant so that the above relation simplifies to
@@
Generalizing our schemes to a general initial velocity $V_{0}$ is straightforward.
For a finite element discretization we closely follow [15]. Let $\mathcal{S}^{m}_{h}$ be a polyhedral surface, i.e. a union of nondegenerate triangles with no hanging vertices (see [10, p. 164]), approximating the closed surface $\mathcal{S}(t_{m})$. Let $\mathcal{S}^{m}_{h}=\bigcup_{j=1}^{J}\overline{\sigma^{m}_{j}}$, where $\{\sigma^{m}_{j}\}_{j=1}^{J}$ is a family of mutually disjoint open triangles with vertices $\{p^{m}_{k}\}_{k=1}^{K}$ and set $h_{m}:=\max_{j=1\to J}{\rm diam}(\sigma^{m}_{j})$. Then let
@@
while in the case $g(s)=1+\tfrac{1}{2}s$
@@ 3.2 The axisymmetric case
$\kappa\nu-\frac{\nu\cdot e_{1}}{x\cdot e_{1}}\nu=\frac{1}{|x_{\rho}|}\bigl{(}\frac{x_{\rho}}{|x_{\rho}|}\bigr{)}_{\rho}-\frac{1}{|x_{\rho}|}\frac{x_{\rho}\cdot e_{2}}{x\cdot e_{1}}\tau^{\perp}\quad\mbox{ in }I\times(0,T],$ (3.16)
@@
$=\widehat{g}^{m}_{j}\left[\frac{1}{2q^{m}_{j+1}}\delta^{-}x^{m+1}_{j+1}-\frac{1}{2q^{m}_{j}}\delta^{-}x^{m+1}_{j}+\frac{1}{2q^{m}_{j+1}}\delta^{-}x^{m-1}_{j+1}-\frac{1}{2q^{m}_{j}}\delta^{-}x^{m-1}_{j}-\frac{\delta^{1}x^{m}_{j}\cdot e_{2}}{x^{m}_{j}\cdot e_{1}}(\theta^{m}_{j})^{\perp}\right]$
@@
$x^{m+1}_{0}\cdot e_{1}=0;\quad\delta^{+}x^{m+1}_{0}\cdot e_{2}=\tfrac{1}{4}h\frac{(q^{m}_{1})^{2}}{\widehat{g}^{m}_{0}}\frac{x^{m+1}_{0}-2x^{m}_{0}+x^{m-1}_{0}}{(\Delta t)^{2}}\cdot e_{2},$ (3.17b)
$x^{m+1}_{J}\cdot e_{1}=0;\quad\delta^{-}x^{m+1}_{J}\cdot e_{2}=-\tfrac{1}{4}h\frac{(q^{m}_{J})^{2}}{\widehat{g}^{m}_{J}}\frac{x^{m+1}_{J}-2x^{m}_{J}+x^{m-1}_{J}}{(\Delta t)^{2}}\cdot e_{2},$ (3.17c)
where $\widehat{g}^{m}_{j}=g(|\frac{x^{m}_{j}-x^{m-1}_{j}}{\Delta t}|^{2})$. In order to motivate the second conditions in (3.17b) and (3.17c), we first calculate with the help of a Taylor expansion and (2.8b) that
$\frac{x(h,t_{m+1})-x(0,t_{m+1})}{h}\cdot e_{2}\approx x_{\rho}(0,t_{m+1})\cdot e_{2}+\tfrac{1}{2}hx_{\rho\rho}(0,t_{m})\cdot e_{2}=\tfrac{1}{2}hx_{\rho\rho}(0,t_{m})\cdot e_{2}.$ (3.18)
Moreover, we deduce from (3.16), [11, (A3)] and the identity $x_{\rho}(0,t_{m})=|x_{\rho}(0,t_{m})|e_{1}$ that
$\lim_{\rho\searrow 0}\bigl{(}\kappa\nu-\frac{\nu\cdot e_{1}}{x\cdot e_{1}}\nu\bigr{)}(\rho,t_{m})$ $=\left[\frac{x_{\rho\rho}}{|x_{\rho}|^{2}}-\frac{x_{\rho\rho}\cdot x_{\rho}}{|x_{\rho}|^{4}}x_{\rho}\right](0,t_{m})-\lim_{\rho\searrow 0}\left[\frac{1}{|x_{\rho}|^{2}}\frac{x_{\rho}\cdot e_{2}}{x\cdot e_{1}}x_{\rho}^{\perp}\right](\rho,t_{m})$
$=2\frac{x_{\rho\rho}(0,t_{m})\cdot e_{2}}{|x_{\rho}(0,t_{m})|^{2}}e_{2}.$
In addition, we see from (2.8b) that $x_{t}(0,t_{m})\cdot e_{1}=0$ and $\tau_{t}(0,t_{m})\cdot e_{2}=0$. Hence $(x_{t}\cdot\tau_{t})(0,t_{m})=0$, so that (2.8a) yields
$x_{tt}(0,t_{m})\cdot e_{2}=g(|x_{t}(0,t_{m})|^{2})e_{2}\cdot\lim_{\rho\searrow 0}\bigl{(}\kappa\nu-\frac{\nu\cdot e_{1}}{x\cdot e_{1}}\nu\bigr{)}(\rho,t_{m})=2g(|x_{t}(0,t_{m})|^{2})\frac{x_{\rho\rho}(0,t_{m})\cdot e_{2}}{|x_{\rho}(0,t_{m})|^{2}},$
or equivalently
$\tfrac{1}{2}hx_{\rho\rho}(0,t_{m})\cdot e_{2}=\tfrac{1}{4}h\frac{|x_{\rho}(0,t_{m})|^{2}}{g(|x_{t}(0,t_{m})|^{2})}x_{tt}(0,t_{m})\cdot e_{2}.$
Inserting this relation into (3.18) and replacing derivatives by suitable difference quotients, we are led to (3.17b). A similar argument applies to (3.17c).
@@
$\tfrac{1}{2}(q^{m}_{j}+q^{m}_{j+1})h^{2}\frac{X_{j}}{(\Delta t)^{2}}=\frac{\widehat{g}^{m}_{j}}{2q^{m}_{j+1}}(X_{j+1}-X_{j})-\frac{\widehat{g}^{m}_{j}}{2q^{m}_{j}}(X_{j}-X_{j-1}),\quad j=1,\ldots,J-1,$ (3.19a)
$X_{0}\cdot e_{1}=0;\quad(X_{1}-X_{0})\cdot e_{2}=\tfrac{1}{4}h^{2}\frac{(q^{m}_{1})^{2}}{\widehat{g}^{m}_{0}}\frac{X_{0}}{(\Delta t)^{2}}\cdot e_{2},$ (3.19b)
$X_{J}\cdot e_{1}=0;\quad(X_{J}-X_{J-1})\cdot e_{2}=-\tfrac{1}{4}h^{2}\frac{(q^{m}_{J})^{2}}{\widehat{g}^{m}_{J}}\frac{X_{J}}{(\Delta t)^{2}}\cdot e_{2}.$ (3.19c)
Clearly, (3.19) decouples into two tridiagonal linear systems for $X\cdot e_{1}$ and $X\cdot e_{2}$, which are both strictly diagonally dominant. Hence it follows that $X=0$.
@@
In addition we have from [5, (2.17)] that
@@ 4 Numerical results
We remark that in all our computations for (1.2) and (1.3), respectively, natural discrete analogues of $E(t)=\tfrac{1}{2}\int_{\mathcal{S}(t)}(\mathcal{V}_{\mathcal{S}}^{2}+2)\;{\rm d}{\mathcal{H}}^{2}$ and $\widetilde{E}(t)=\int_{\mathcal{S}(t)}e^{\frac{1}{2}\mathcal{V}_{\mathcal{S}}^{2}}\;{\rm d}{\mathcal{H}}^{2}$ were well preserved, recall (1.2) and (1). In particular, in all our computations the relative differences between these discrete analogues at time $t=0$ and at time $t=T$ are always less than $1.5\%$, even though the computations often come very close to a singularity.
@@ 4.1.1 Radial evolutions for a sphere
Our first set of numerical experiments is for the evolution of a sphere. To validate our schemes (3.1) and (3.17) with $g(s)=1$, we begin with a convergence experiment, generalizing the solution from [12, §5.2] to higher space dimensions. It can be shown that a family of spheres in $\mathbb{R}^{3}$ with radius $r(t)$ and outer normal $\nu(t)$ is a solution to (1.2) with $\mathcal{V}_{\mathcal{S}\mid_{t=0}}=V_{0}\in{\mathbb{R}}$ if
@@
and where we employed the time step size choice $\Delta t=0.25h_{0}$. We note that the observed order of convergence approaches the optimal rate $\mathcal{O}(h^{2}_{0})$.
@@
and where we have chosen $\Delta t=h~{}=\frac{1}{J}$. Also for the scheme (3.17) we obtain an optimal convergence rate.
@@ 4.1.2 Axisymmetric evolutions for nonspherical surfaces
We begin with an experiment for an initial $2:1:1$ cigar, when the initial velocity is chosen as $V_{0}=0$, see Figure 3. Again we observe an excellent agreement between the solutions for our schemes (3.1) and (3.17) with $g(s)=1$. At around $t=0.5$ there seems to be the onset of a singularity. Notice that the first kink in the curvature plot at around time $t=0.43$ indicates a temporary increase in the magnitude of the mean curvature at the axis of rotation. However, this does not appear to lead to a singularity. Instead, the shorter part of the surface eventually becomes flat, leading to a blow-up of curvature where the flat part meets the curved part of the surface. To qualitatively confirm these results, we have repeated this simulation also with finer discretization parameters, and the described behaviour is consistent across the different computations.
@@
A simulation for an initial torus with radii $R=2$ and $r=1$ can be seen in Figure 5 for $V_{0}=0$. Here we see a singularity develop, before the torus can shrink to a circle and vanish, when the outer part of the torus becomes flat. This kind of singularity cannot occur for mean curvature flow, and so is unique to the hyperbolic flow considered here. In particular, we recall from [27, Proposition 3] that for mean curvature flow it is known that there exists a critical value $r_{crit}\in(0,1)$, such that for $0<\frac{r}{R}<r_{crit}$ the solution shrinks to a circle, while for $r_{crit}<\frac{r}{R}<1$ it closes up the hole at the origin. This was also confirmed by numerical experiments in e.g. [2, Figs. 5, 6], [6, Figs. 2, 3] and [1, Figs. 2, 3, 4], where the latter also provides an estimate for $r_{crit}$ of about $0.6415$.
@@
We end this section with a simulation for hyperbolic mean curvature flow for a surface that is the boundary of a nonconvex domain. In particular, as the initial discrete surface $\mathcal{S}^{0}_{h}$ we choose the biconcave disk obtained in [4, Fig. 8]. The total dimensions of $\mathcal{S}^{0}_{h}$ are about $4.2\times 4.2\times 1.1$, and the discrete surface is made up of 3072 elements, with the largest element being about three times the size of the smallest one. The numerical results for the scheme (3.1) are presented in Figure 7. We notice that at the final time $t=T$ the mesh quality has deteriorated significantly, with the largest element now being nearly 500 times the size of the smallest one.
@@
In order to qualitatively confirm the evolution shown in Figure 7, we compute a very similar evolution for the axisymmetric finite difference scheme (3.17) with $g(s)=1$. Here we choose as initial data $x^{0}$ the curve displayed in the middle of [8, Fig. 5]. The obtained simulation is shown in Figure 8, which agrees very well with the full 3d results in Figure 7. In particular, at the side of the disk a concavity appears, which then leads to a blow-up in curvature and a resulting singularity for the flow.